OpenStudy (anonymous):

how does this problem differ from a regular synthetic division problem? No divisor is given. P(-3) for P(x)=x^4-2x^3-4x+4

OpenStudy (anonymous):

the difference here is you just plug in x value to P(x), while in synthetic div you are dividing or using addition or subtractions of terms

OpenStudy (anonymous):

i know how to do syn. divi but don't know how to set up this problem

OpenStudy (anonymous):

P(x)=x^4-2x^3-4x+4 P(-3)=(-3)^4-2(-3)^3-4(-3)+4 P(-3)=151 youl arrive with the same answer on doing synthetic div

OpenStudy (anonymous):

I did it that way and did get that answer, but now how do I do it with syn. div since I don't have a divisor

OpenStudy (anonymous):

its -3

OpenStudy (anonymous):

but I don't get 151 are my numbers 1 -2 -4 4?

OpenStudy (anonymous):

try 1 -2 0 -4 4

OpenStudy (anonymous):

I did put in ) and I got -119

OpenStudy (anonymous):

i did put in 0 and got -119

OpenStudy (anonymous):

it should have a coefficient of ox^2 w/c is missing there right?

OpenStudy (anonymous):

right there is no x^2

OpenStudy (anonymous):

try it again youl arrive at a remainder of R=151l

OpenStudy (anonymous):

got it a dumb math mistake thanks

OpenStudy (anonymous):

the answer is 1 -5 15 -49 R151 its now (x+3)(x^3 -5x^2 +15x -49) + R=151

OpenStudy (anonymous):

ok good luck now..YW...have fun...:D